If Marie works 14 hours and tom works 17 hours doing the same job. How long would it itake them doing the same job?
Hint Rate X Time + amount of work

Answer by ptaylor(2198)

I assume the problem wants to know how long it takes them working together to do the job??
Let x=amount of time it takes them working together
Then together they work at the rate of 1/x job per hour
Marie works at the rate of 1/14 job per hour
Tom works at the rate of 1/17 job per hour
so, our equation to solve is:
1/14 + 1/17= 1/x multiply each term by 238x
17x+14x=238
31x=238
x=7.677 hours---------------------------time it takes them working together
CK
(1/14)*(238/31)+ (1/17)*(238/31)=1 (1 job, that is)
17/31 + 14/31=1
31/31=1
1=1

Hope this helps------ptaylor